A 1 kg-body moves upwards across the vertical wall due to the force F=20 N applied at the angle a = 30 with vertical. Find the acceleration of the body if the coefficient of friction is k = 0.05
Answer #1

here ,

let the acceleration is a

Normal force , N = 20 * sin(30) = 10 N

Now, Using second law of motion

20 * cos(30) - 0.05 * 10 - 1 * 10 = 1 * a

a = 6.82 m/s^2

the acceleration due to body is 6.82 m/s^2